# Outlier Gradient Analysis: Efficiently Improving Deep Learning Model Performance via Hessian-Free Influence Functions

### v2

@@ Abstract
Influence functions offer a robust framework for assessing the impact of each training data sample on model predictions, serving as a prominent tool in data-centric learning. Despite their widespread use in various tasks, the strong convexity assumption on the model and the computational cost associated with calculating the inverse of the Hessian matrix pose constraints, particularly when analyzing large deep models. This paper focuses on a classical data-centric scenario—trimming detrimental samples—and addresses both challenges within a unified framework. Specifically, we establish an equivalence transformation between identifying detrimental training samples via influence functions and outlier gradient detection. This transformation not only presents a straightforward and Hessian-free formulation but also provides profound insights into the role of the gradient in sample impact. Moreover, it relaxes the convexity assumption of influence functions, extending their applicability to non-convex deep models. Through systematic empirical evaluations, we first validate the correctness of our proposed outlier gradient analysis on synthetic datasets and then demonstrate its effectiveness in detecting mislabeled samples in vision models, selecting data samples for improving performance of transformer models for natural language processing, and identifying influential samples for fine-tuned Large Language Models.
@@ 1 Introduction
Data-centric learning focuses on enhancing algorithmic performance from the perspective of the training data (Oala et al., 2023). In contrast to model-centric learning, which designs novel algorithms or optimization techniques for performance improvement with fixed training data, data-centric learning operates with a fixed learning algorithm while modifying the training data through trimming, augmenting, or other methods aligned with improving utility (Zha et al., 2023). Data-centric learning holds significant potential in many areas such as model interpretation, subset training set selection, data generation, noisy label detection, active learning, and others (Chhabra et al., 2024; Kwon et al., 2024).
@@
The essence of data-centric learning lies in estimating data influence, also known as data valuation (Hammoudeh and Lowd, 2022), in the context of a learning task. Intuitively, the impact of an individual data sample can be measured by assessing the change in learning utility when training with and without that specific sample. This leave-one-out influence (Cook and Weisberg, 1982) provides a rough gauge of the relative data influence of the specific sample on the otherwise full fixed training set. On the other hand, Shapley value (Ghorbani and Zou, 2019; Jia et al., 2019), originating from cooperative game theory, quantifies the increase in value when a group of samples collaborates to achieve the learning goal. Unlike leave-one-out influence, Shapley value represents the weighted average utility change resulting from adding the point to different training subsets. Despite the absence of assumptions on the learning model, the aforementioned retraining-based methods incur significant computational costs, especially for large-scale data analysis and deep models (Hammoudeh and Lowd, 2022).
@@
Diverging from the retraining-based category of approaches, influence functions assess data influence without requiring model retraining (Koh and Liang, 2017). They measure the effect of changing an infinitesimal weight of training samples based on a utility-evaluating function. While influence functions can be accurate or acceptable proxies for convex and certain shallow models (Koh and Liang, 2017), their applicability to deep models is constrained by the strong convexity assumption and the computational cost linked to calculating the inverse of the Hessian matrix (Basu et al., 2020a).
Our Contributions. In this paper, we delve into a classical data-centric scenario: trimming detrimental samples, also known as best subset selection (Bertsimas et al., 2016). We strive to tackle two significant challenges associated with influence functions within a unified framework. Our contributions are as follows:
@@
  • We build an equivalence transformation between identifying detrimental training samples via influence functions and outlier detection on the gradient space of samples, and propose our outlier gradient analysis approach. The theoretical transformation features a straightforward and Hessian-free formulation, and reduces the computational cost associated with the Hessian matrix and its inverse.
  • Furthermore, our transformation relaxes the convex assumption of influence functions and extends their applicability from simple convex learning models to non-convex deep models via the Taylor expansion to high-order gradients. We utilize the first-order gradient for outlier analysis. Notably, it can be obtained during model training, and no additional high-order gradients are necessary for our outlier gradient analysis approach.
  • Empirically we demonstrate that outlier gradient analysis can lead to performance improvements compared to existing influence-based approaches in deep learning while being computationally more efficient. We first demonstrate the efficacy of our framework compared to baselines on synthetic datasets. Subsequently, we showcase the usefulness of outlier gradient analysis on trimming mislabeled samples in vision datasets across a variety of noisy regimes. We also consider Natural Language Processing (NLP) applications by conducting experiments on data selection for fine-tuning deep transformer models such as RoBERTa Liu et al. (2019) and influential data identification for text generation tasks on fine-tuned Large Language Models (LLMs). As our results show, outlier gradient analysis obtains performance gains in a computationally efficient manner across diverse tasks, models, and application paradigms.
@@ 2 Related Work
Retraining-Based Influence Estimation. Influence estimation and data valuation approaches can be generally categorized as either retraining-based or gradient-based (Hammoudeh and Lowd, 2022). Retraining-based methods consist of the classical leave-one-out influence approach (Cook and Weisberg, 1982) which consists of removing one (or a group of) training sample(s) at a time, and retraining the model to measure sample influence via performance change. While useful as an ideal baseline, on deep models this is clearly an untenable influence approach as it necessitates computationally expensive model retrainings every time a sample is removed. Other representative methods include Shapley value approaches (Ghorbani and Zou, 2019; Jia et al., 2019; Kwon and Zou, 2022), which are model agnostic, but also computationally untenable for large datasets and deep models due to exponential time complexity. Computationally efficient approaches such as KNN-Shap (Jia et al., 2018) tend to make assumptions about the learning model and hence are not directly applicable to the (deep) models being considered. It is evident that for deep learning influence estimation, retraining-based approaches are not ideal, and hence, gradient-based methods are more suitable.
@@
Gradient-Based Influence Estimation. For models trained using gradient descent, gradient-based influence approaches can be used to approximately estimate influence without requiring retraining. The seminal work in this category is that of Koh and Liang (2017), which utilizes a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. However, the limiting assumption is that the model and loss function are convex, which is not true for deep models. Follow-up works such as representer point (Yeh et al., 2018) and Hydra (Chen et al., 2021) inherit these convexity assumptions and suffer from similar issues of applicability. Moreover, most modern influence function approaches (Schioppa et al., 2022; Feldman and Zhang, 2020) tend to be too computationally expensive for large models, and cannot run in reasonable time. More recently, Kwon et al. (2024) proposed DataInf, which efficiently computes influence even for large models simply be replacing the inverse Hessian computation with an easily computable closed-form expression, but their framework implicitly assumes convexity as well. Some approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019). Other work in this domain has investigated traditional Hessian-based influence functions for relabeling samples (Kong et al., 2021), and how they can be scaled to use with LLMs (Grosse et al., 2023). Recent work has also investigated the role of the Hessian and convexity in influence estimation (Schioppa et al., 2024). In contrast, our work aims to circumvent these issues by operating on the gradient space in a skillful manner, and obviates the need for assuming model convexity. Hence, our work paves the way for an efficient and accurate influence estimation framework and adds to the “influence function toolset” for deep models and large datasets. Finally, recent work has also found that self-influence (influence computed on training samples) can be beneficial in detecting detrimental samples (Bejan et al., 2023; Thakkar et al., 2023) so we compare our proposed method with these approaches as well.
Miscellaneous Data-Centric Learning. Many works in the data-centric learning domain study other relevant research questions. For instance, datamodels (Ilyas et al., 2022) also estimate training sample contributions, but only for one test sample at a time. Data efficiency approaches (Jain et al., 2023; Paul et al., 2021; Killamsetty et al., 2021) aim to accelerate deep learning training time via subset selection. Data pruning approaches based on novel approximations for leave-one-out influence estimation (Tan et al., 2024) and the model’s generalization gap (Yang et al., 2022) have also been proposed. Model pruning via generalized influence functions has also been studied in (Lyu et al., 2023). Antidote data augmentation (Chhabra et al., 2022; Li et al., 2023) methods aim to generate synthetic data samples to improve model performance, whereas feature selection approaches (Hall, 1999; Cai et al., 2018) seek to optimize the feature space to only those important for model performance. Active learning (Cohn et al., 1996) methods aim to iteratively identify optimal samples to annotate given a large unlabeled training data pool (Liu et al., 2021; Nguyen et al., 2022; Wei et al., 2015). Finally, works on poisoning attacks seek to analyze model robustness by perturbing training set samples (Solans et al., 2021; Mehrabi et al., 2021; Chhabra et al., 2023) under natural input constraints. The study of training sample influence has also been extended to recent generative models, such as diffusion models (Dai and Gifford, 2023), through the use of ensembles.
@@ 3 Proposed Approach
We first introduce influence functions conceptually and outline their classical application in identifying detrimental samples. We then detail our equivalence transformation by converting the research question into a gradient space outlier analysis problem. Subsequently, we provide insights for extending influence functions to non-convex learning models and propose our outlier gradient analysis approach.
@@ 3.1 Preliminaries on Influence Functions
One key application of influence functions lies in identifying detrimental samples. An intuitive approach involves training the model both with and without the specific training sample and assessing the resulting changes using metrics like training/validation loss. In other words, if the performance improves when excluding a particular sample, it is deemed detrimental to the learning task. By computing the influence score without needing to retrain the model, one can estimate the impact of a sample to assess if it is beneficial or detrimental, as follows:
@@
~$\tilde{\mathcal{I}}(z_{j})$ can be regarded as a the discrete version of $\mathcal{I}(z_{j})$. Specifically, a positive value for $\mathcal{I}(z_{j})$ means that removing the sample $z_{j}$ enhances the model’s utility, and that $z_{j}$ is a detrimental sample.
While influence functions offer a swift estimation of the training sample effect without the need for costly model retraining, their practical applications are constrained by two prominent drawbacks. The initial limitation lies in the necessity of a strictly convex loss function to guarantee the existence of the inverse of the Hessian matrix. The second challenge pertains to the considerable computational expense associated with calculating the inverse of the Hessian matrix. For the first challenge, several possible solutions exist: (1) a linear model can be used as a surrogate on the embeddings obtained via the non-convex model (Chhabra et al., 2024); (2) a damping term can be added to the non-convex model such that its Hessian becomes positive definite and invertible (Han et al., 2020); and (3) for certain tasks specific or second-order influence functions can be derived (Basu et al., 2020b; Alaa and Van Der Schaar, 2020). For the second challenge, various matrix inverse techniques are employed to expedite the computation process, including LiSSA optimization (Koh and Liang, 2017) and swapping the order of the matrix inversion (Kwon et al., 2024). Considerable efforts have been dedicated to addressing the aforementioned challenges with promising results; this paper aims to unify and confront both challenges under a single framework. We endeavor to present a straightforward and elegant solution for extending influence functions to non-convex loss functions.
@@ 3.2 Bridging Influence Estimation and Outlier Analysis
Now we transform the problem of identifying detrimental samples via influence estimation to an outlier analysis problem in the gradient space. Upon scrutinizing the influence estimation of $z_{j}$ in Eq. (1), it becomes evident that the influence score is the result of three terms, with the first two being constant across all training samples and independent of $z_{j}$. While all three terms contribute to the concrete value of the influence score, it is the final term $\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})$ that assumes a decisive role in determining whether $z_{j}$ is a beneficial or detrimental sample. This is because only the third term has $z_{j}$ as input. With the following observation, we can build the connection between identifying detrimental samples via influence estimation and outlier analysis.
@@
The majority of training samples positively contribute to the model’s utility, with only a small subset (with respect to the overall size of the training set) exhibiting detrimental effects.
Clearly, Observation 3.1 holds true as the empirical loss is an average of error between predictive and true values over all training samples. Hence, detrimental samples can be regarded as outliers compared to the beneficial sample majority. Based on Observation 3.1 and the decisive role of $\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})$ in influence estimation, we have the following proposition:
Proposition 3.2.
There exists an outlier analysis algorithm for detecting outliers in the gradient space of samples, such that determining whether a training sample beneficially or detrimentally impacts model utility through influence estimation is equivalent to applying this outlier analysis algorithm on the gradient space.
Proposition 3.2 establishes a conceptual equivalence transformation between the identification of detrimental training samples via influence estimation and the detection of outliers in the gradient space. This theoretical transformation not only features a straightforward and Hessian-free formulation, reducing the computational cost associated with the Hessian matrix and its inverse, but also yields profound insights into the role of the gradient in sample impact beyond model optimization. Furthermore, this transformation relaxes the convexity assumption of influence functions. While the influence estimation in Eq. (1) is a first-order Taylor expansion, sufficient for convex loss functions, non-convex loss functions demand higher-order Taylor expansions for accurate estimation (Basu et al., 2020b; Alaa and Van Der Schaar, 2020). The higher-order Taylor expansion of influence estimation is the summation of the product of three terms with different orders, where the first term $\sum_{z\in T/V}\nabla_{\hat{\theta}}\ell(z;\hat{\theta})^{\top}$ is a constant. In a similar vein, our transformation for high-order Taylor extensions renders all Hessian-related terms, including their high-order formulations and inverses, constants as well. This characteristic alleviates the need for gradients with different orders in determining whether a sample is beneficial or detrimental, even for non-convex models.
@@ 3.3 Our Approach: Outlier Gradient Analysis
As demonstrated in Proposition 3.2, outlier analysis can effectively evaluate the discrete influence of training samples. Notably, we can circumvent the need for computing and inverting the Hessian and adapt influence function to non-convex deep models while still measuring discrete influence via Eq. (2). The primary contribution and discovery of our work lies in the realization that simple and efficient outlier analysis techniques can be applied to the gradient space for a discrete estimation of which samples are beneficial or detrimental to the model’s utility.
As Proposition 3.2 does not prescribe a specific outlier detection algorithm, one of our choices for outlier analysis is the Isolation Forest (iForest) (Liu et al., 2008) algorithm due to several factors. Firstly, iForest boasts a linear time complexity with a low constant, requiring minimal memory, rendering it well-suited for handling the high-dimensional gradient space inherent in deep models. Secondly, iForest constructs an ensemble of iTrees, where each iTree builds partial models and employs sub-sampling, demonstrating the ability to identify a suitable subspace for the detection of detrimental samples. Thirdly, iForest is known for its simplicity and effectiveness in identifying outliers that are non-linearly separated from inliers. Along with iForest, we also consider two simple outlier analysis approaches based on L1-norm and L2-norm thresholding, that work well in practice (Knorr et al., 2000).
@@
4 Correctness Verification on Synthetic Data
We seek to validate the correctness of our proposed idea and showcase the effectiveness of our outlier gradient analysis method on two synthetic 2D toy datasets [^1] Comprehensive details regarding datasets and model training for experiments are provided in Appendix A. and two models for binary classification in Figure 1. In this figure, subfigures A-D present a linear dataset employing a Logistic Regression model, while subfigures E-H exhibit a non-linear dataset utilizing a non-convex Multilayer Perceptron (MLP) model as the base model. Specifically, subfigures A and B depict the training and test sets of a linearly separable dataset comprising 150 and 100 samples, respectively. Notably, the training set includes 10 manually generated noisy samples with misspecified labels. Subfigure C displays the influence score of each training sample, computed using Eq. (1), and subfigure D provides a visualization of the gradient space. Similarly, subfigures E and F represent the training and test sets of the two half moons dataset, with the training set consisting of 250 samples and the test set of 100 samples, equally distributed between two classes. The training set in this case also contains 20 noisy samples. Subfigures G and H showcase the influence score and gradient space of the non-convex case.
Figure 1: Illustrating our outlier gradient analysis approach on two synthetic datasets and convex/non-convex models. Subfigures A-D showcase our outlier gradient analysis approach on a 2D linearly separable synthetic dataset. This dataset includes a small subset of detrimental samples with incorrect labels used to train a Logistic Regression binary classification model. Meanwhile, subfigures E-H depict our outlier gradient analysis on a non-linear synthetic dataset with mislabeled samples employed in training a Multilayer Perceptron (MLP) neural network. In subfigures A and E, the training sets are represented with class labels 0 (red) and 1 (blue) in the convex and non-convex cases, respectively. Detrimental samples with incorrect class labels are marked with $\times$, while regular samples are marked with $\circ$. Subfigures B and F denote the test sets used to evaluate model performance in the convex and non-convex scenarios. Subfigures C and G display the influence scores calculated by Eq. (1). Upon running our outlier analysis algorithm on the gradient space of the linear convex model (subfigure C), most detrimental samples are identified (depicted in orange), while regular points are considered inliers. Notably, these outliers exhibit lower influence scores, as influence functions prove accurate for convex models. In contrast, subfigure G demonstrates that influence scores are not reliable indicators for detecting detrimental samples in the non-convex case. However, after applying outlier analysis on the gradient space of the non-convex MLP model, most detrimental samples are detected (depicted in orange). Subfigures D and H showcase the gradient space obtained for each sample from the Logistic Regression and MLP models, respectively. It is evident that the outlier samples correspond to detrimental samples with mislabeled classes, which is linearly or non-linearly separated from inliers. Note that the benefits of outlier gradient trimming can be clearly observed—removing predicted outlier samples via iForest and retraining the MLP enhances classification performance from $90\%\rightarrow 96\%$ on the test set (refer to Table 1).
In the linear case, as illustrated in subfigure C, the influence score proves to be a reliable indicator for distinguishing detrimental samples from beneficial ones. Notably, detrimental samples exhibit large negative scores, while other samples display positive or nearly zero values. Additionally, subfigure D affirms that these detrimental samples are distinctly separated in the gradient space, confirming the validity of the equivalent transformation outlined in Proposition 3.2. However, the limitations of influence scores become evident in the context of non-convex models, as observed in subfigure G, where the influence scores of detrimental samples are mixed with those of normal ones. Nevertheless, in the gradient space illustrated in subfigure H, the detrimental samples are effectively isolated from inliers. Notably, our method does not rely on the Hessian for computing influence; instead, it operates directly on the gradient space using outlier analysis techniques. This highlights that our gradient outlier analysis approach effectively relaxes the convexity assumption associated with influence functions.
Table 1: Outlier detection and classification performance of noisy label correction and influence-based approaches including our proposed outlier gradient trimming on the two half moons dataset (top-2 performers in bold).

### v3

# Outlier Gradient Analysis: Efficiently Identifying Detrimental Training Samples for Deep Learning Models

@@ Abstract
A core data-centric learning challenge is the identification of training samples that are detrimental to model performance. Influence functions serve as a prominent tool for this task and offer a robust framework for assessing training data influence on model predictions. Despite their widespread use, their high computational cost associated with calculating the inverse of the Hessian matrix pose constraints, particularly when analyzing large-sized deep models. In this paper, we establish a bridge between identifying detrimental training samples via influence functions and outlier gradient detection. This transformation not only presents a straightforward and Hessian-free formulation but also provides insights into the role of the gradient in sample impact. Through systematic empirical evaluations, we first validate the hypothesis of our proposed outlier gradient analysis approach on synthetic datasets. We then demonstrate its effectiveness in detecting mislabeled samples in vision models and selecting data samples for improving performance of natural language processing transformer models. We also extend its use to influential sample identification for fine-tuning Large Language Models.
@@ 1 Introduction
Data-centric learning focuses on enhancing algorithmic performance from the perspective of the training data (Oala et al., 2023). In contrast to model-centric learning, which designs novel algorithms or optimization techniques for performance improvement with fixed training data, data-centric learning operates with a fixed learning algorithm while modifying the training data through trimming, augmenting, or other processing for improving utility (Zha et al., 2023). Data-centric learning holds significant potential in many areas such as model interpretation, subset training set selection, data generation, noisy label detection, active learning, and others (Chhabra et al., 2024; Kwon et al., 2024).
@@
The essence of data-centric learning lies in estimating data influence, also known as data valuation in the context of a learning task (Hammoudeh and Lowd, 2022),. Intuitively, the impact of an individual data sample can be measured by assessing the change in learning utility when training with and without that specific sample. This leave-one-out influence (Cook and Weisberg, 1982) provides a rough gauge of the relative data influence of the specific sample on the otherwise full fixed training set. Shapley value (Ghorbani and Zou, 2019; Jia et al., 2019), originating from cooperative game theory, quantifies the increase in value when a group of samples collaborates to achieve the learning goal. Unlike leave-one-out influence, Shapley value represents the weighted average utility change resulting from adding the sample to different training subsets. Despite the absence of assumptions on the learning model, the aforementioned retraining-based methods incur significant computational costs, especially for large-scale data analysis and deep models (Schioppa et al., 2022).
@@
A popular choice for data valuation applications, such as identifying training samples detrimental to model performance, are influence functions (Koh and Liang, 2017). Essentially, influence functions assess data influence without requiring model retraining. They measure the effect of changing an infinitesimal weight of training samples based on a utility-evaluating function. While influence functions can be accurate or acceptable proxies for convex and certain shallow models, their applicability to deep models is constrained by the strong convexity assumption and the computational cost linked to calculating the inverse of the Hessian matrix (Basu et al., 2020a).
Our Contributions. In this paper, we delve into the classical data-centric problem: identifying/trimming detrimental samples. We tackle the computational challenge of the inverse of the Hessian matrix in influence functions in the context of detrimental sample identification and removal. Our major contributions are as follows:
@@
  • We build a bridge between identifying detrimental training samples via influence functions and outlier detection on the gradient space of samples, and propose our outlier gradient analysis approach. The transformation features a straightforward and Hessian-free formulation, and reduces the computational cost associated with the Hessian matrix and its inverse.
  • Empirically, we utilize both linear and non-linear synthetic datasets to illustrate the ineffectiveness of the current Hessian approximation and to validate our hypothesis regarding outlier gradient analysis, showcasing our method’s high accuracy in identifying mislabeled detrimental samples.
  • Subsequently, we demonstrate the effectiveness of outlier gradient analysis in trimming mislabeled samples from vision datasets across various noise regimes. Additionally, we explore textual applications by conducting experiments on data selection for fine-tuning deep transformer models and identifying influential data for text generation tasks using fine-tuned Large Language Models..
@@ 2 Related Work
Retraining-Based Influence Estimation. Influence estimation approaches can be generally categorized as either retraining-based or gradient-based (Hammoudeh and Lowd, 2022). Retraining-based methods consist of the classical leave-one-out influence approach (Cook and Weisberg, 1982), which consists of removing one training sample at a time, and retraining the model to measure sample influence via performance change. Other representative methods include Shapley value approaches (Ghorbani and Zou, 2019; Jia et al., 2019; Kwon and Zou, 2022), which are model agnostic, but also computationally untenable for large datasets and deep models due to exponential time complexity. Computationally efficient approaches such as KNN-Shap (Jia et al., 2018) can only employ KNN classifiers and hence are not directly applicable to the deep models.
@@
Gradient-Based Influence Estimation. For models trained using gradient descent, gradient-based influence approaches can be used to approximately estimate influence without requiring retraining. The seminal work in this category is that of Koh and Liang (2017), which utilizes a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. However, the limiting underlying assumption in the formulation is that the model and loss function are convex, which is not true for deep models. Follow-up works such as representer point (Yeh et al., 2018) and Hydra (Chen et al., 2021) inherit these convexity assumptions and suffer from similar issues of applicability. While influence functions have been used for numerous applications in data-centric learning (Feldman and Zhang, 2020; Chhabra et al., 2024; Richardson et al., 2023), they tend to be too computationally expensive for large models, and cannot run in reasonable time. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2024), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for large models. Some approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also investigated the role of the Hessian and convexity in influence estimation (Schioppa et al., 2024). In contrast, our work aims to circumvent these issues for detrimental sample identification by operating on the gradient space in a skillful manner. Hence, our work paves the way for an efficient and accurate detrimental sample identification framework and adds to the “influence function toolset” for deep models and large datasets. Finally, recent work has also found that self-influence (influence computed on training samples) can be beneficial in detecting detrimental samples (Bejan et al., 2023; Thakkar et al., 2023). For related works on miscellaneous data-centric learning, please refer to Appendix A.
@@ 3 Proposed Approach
We first introduce influence functions conceptually and outline how they are applied to the task of detrimental samples identification. We then detail our transformation by converting the original formulation into a gradient space outlier analysis problem. Subsequently, we provide insights for extending influence functions to non-convex learning models and propose our outlier gradient analysis approach.
@@ 3.1 Preliminaries on Influence Functions
One key application of influence functions lies in identifying detrimental samples. This is because an intuitive way of assessing whether a sample is detrimental is by training the model both with and without the specific training sample and computing metrics like training/validation loss. In other words, if the performance improves when excluding a particular sample, it is deemed detrimental to the learning task. By computing the influence score without needing to retrain the model, one can estimate the impact of a sample to assess if it is beneficial or detrimental, as follows:
@@
~$\tilde{\mathcal{I}}(z_{j})$ can be regarded as a the discrete version of $\mathcal{I}(z_{j})$. Specifically, a value of 0 for ~$\tilde{\mathcal{I}}(z_{j})$ means that removing the sample $z_{j}$ enhances the model’s utility, and that $z_{j}$ is a detrimental sample.
Remark. While influence functions offer a swift estimation for identifying detrimental training samples without the need for costly model retraining, their practical applications to large models are constrained by two prominent drawbacks. The first limitation lies in the necessity of a strictly convex loss function to guarantee the existence of the inverse of the Hessian matrix. The second challenge pertains to the considerable computational expense associated with calculating the inverse of the Hessian. For the first challenge, several possible solutions have been proposed: (1) a convex surrogate model can be used instead of the non-convex model (Chhabra et al., 2024); (2) a damping term can be added to the Hessian to ensure it is positive definite and invertible (Han et al., 2020); and (3) alternative formulations (Basu et al., 2020b; Alaa and Van Der Schaar, 2020) can be used (e.g. the Gauss Newton Hessian (Grosse et al., 2023) instead of the standard Hessian). Note that some studies bypass the convexity assumption and directly apply influence functions to deep models, yielding effective results. (Grosse et al., 2023). For the second challenge, various matrix inverse techniques are employed to expedite the computation process, including LiSSA optimization (Koh and Liang, 2017) and swapping the order of the matrix inversion (Kwon et al., 2024), among several others. Considerable efforts have been dedicated to addressing the aforementioned challenges with promising results– however, in this paper we target the second challenge for identifying/removing detrimental samples.
@@ 3.2 Bridging Influence Estimation and Outlier Analysis
We transform the problem of identifying detrimental samples via influence estimation to an outlier analysis problem in the gradient space. Upon scrutinizing the influence estimation of $z_{j}$ in Eq. (1), it becomes evident that the influence score is the result of three terms, with the first two remaining the same across all training samples and not solely dependent on $z_{j}$. While all three terms contribute to the concrete value of the influence score, it is the final term $\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})$ that assumes a decisive role in determining whether $z_{j}$ is a beneficial or detrimental sample. This is because the third term has $z_{j}$ as the only training sample as an input. With the following observation below regarding detrimental samples, we can build the connection between identifying detrimental samples via influence estimation and outlier analysis:
@@
For a converged model trained using empirical risk minimization, the majority of training samples positively contribute to the model’s utility, and a much smaller subset than beneficial samples (with respect to the overall size of the training set) exhibits detrimental effects.
Clearly, Observation 3.1 holds true as the empirical loss is an average of error between predictive and true values over all training samples. Hence, detrimental samples can be regarded as a minority outlier set compared to the beneficial sample majority. Based on Observation 3.1 and the decisive role of $\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})$ in influence estimation, we have the following hypothesis:
Hypothesis 3.2.
There exist outlier analysis algorithms capable of detecting detrimental samples in the gradient space. This algorithm would enable us to evaluate whether a training sample positively or negatively impacts model utility through influence estimation, effectively equating this evaluation with the application of the outlier analysis algorithm in the gradient space.
Hypothesis 3.2 establishes a conceptual transformation between the identification of detrimental training samples via influence estimation and the detection of outliers in the gradient space. The outlying nature of detrimental samples has also been observed in past work (Kim et al., 2024). This transformation not only features a straightforward and Hessian-free formulation, reducing the computational cost associated with the Hessian matrix and its inverse, but also yields insights into the role of the gradient in sample impact beyond model optimization.
@@ 3.3 Our Approach: Outlier Gradient Analysis
As demonstrated in Hypothesis 3.2, outlier analysis can effectively be used to evaluate the discrete influence of training samples. Notably, we can circumvent the need for computing and inverting the Hessian for non-convex deep models by measuring discrete influence via Eq. (2). The primary contribution and discovery of our work lies in the realization that simple and efficient outlier analysis techniques can be applied to the gradient space for a discrete estimation of which samples are beneficial or detrimental to the model’s utility.
@@
As Hypothesis 3.2 cannot prescribe a specific outlier detection algorithm, one of our choices for outlier analysis is the Isolation Forest (iForest) algorithm (Liu et al., 2008), owing to several factors. Firstly, iForest boasts a linear time complexity with a low constant, requiring minimal memory, rendering it well-suited for handling the high-dimensional gradient space inherent in deep models. Secondly, iForest constructs an ensemble of iTrees, where each iTree builds partial models and employs sub-sampling, demonstrating the ability to identify a suitable subspace for the detection of detrimental samples. Thirdly, iForest is known for its simplicity and effectiveness in identifying outliers that are non-linearly separated from inliers. Along with iForest, we also consider two simple outlier analysis approaches based on L1-norm and L2-norm thresholding, that work well in practice (Knorr et al., 2000).
@@
4 Hypothesis Verification on Synthetic Data
Table 1: Outlier detection and classification performance of noisy label correction and influence-based approaches including our proposed outlier gradient trimming on the two half moons dataset (top performer in bold).